# Recent Advances in Discrete Speech Tokens: A Review

### v1

@@ IV-F Speech Token Vocoders
Polyak et al. [152] first explores speech resynthesis from discrete speech units by a HifiGAN [179] augmented with discretized pitch units and speaker embedding inputs. The vec2wav model in VQTTS [136, 180, 181, 182] improves this vocoder by a Conformer [92] frontend module before HifiGAN generator. The CTX-vec2wav vocoder in UniCATS [119] proposes a position-agnostic cross-attention mechanism that effectively integrates timbre information from surrounding acoustic contexts without the need for pretrained speaker embeddings. This makes it more timbre-controllable and suitable for zero-shot TTS and VC [120]. Its descendant vec2wav 2.0 [183] further advances the timbre controllability by SSL timbre features and adaptive Snake activations, which is demonstrated by a strong VC performance.
It is also feasible to apply diffusion or flow matching algorithms in token vocoders [184, 185, 40]. There, the discrete tokens are treated as a condition for diffusion or flow matching to generate mel-spectrograms, and further converted to waveform by a pretrained mel vocoder. Compared to training a token-to-wav vocoder in an end-to-end fashion, training a token-to-mel model is more convenient and does not need adversarial training. To control timbre better, a mask strategy is introduced into the training process where the model only computes diffusion or flow matching loss on the un-masked part of spectrograms [40]. During inference, spectrogram from speaker prompt conditions the generative process, which can be regarded as a form of “in-context learning”. However, inference efficiency may be compromised for better generation quality with more than one inference steps, and this method is only validated on massive amount of data currently.
@@ V-A Metrics and Existing Benchmarks
  • Signal-level reconstruction metrics: Discrete tokens, especially neural speech codecs, can be evaluated using the same signal-level metrics as in vocoder research. Typical reconstruction metrics include PESQ, STOI, mel distance, STFT distance, GPE, etc.
  • Perceptual reconstruction metrics: Apart from the signal level, there can also be perceptual evaluations of reconstruction performance. This includes intelligibility (often measured by WER, CER, or phone error rates), speaker similarity, subjective or proxy MOS tests, etc.
  • Performance on downstream tasks: Probing tasks can be used to measure the preservation or prominence of certain information in tokens, like ASR, ASV, emotion recognition, and spoken language modeling [176]. Note that this is different from perceptual reconstruction metrics since it operates directly on tokens. Also, as discrete tokens can be applied to more generative speech processing tasks like TTS and VC, it is feasible to compare the performance of these tasks.
  • Semantic/phonetic relevance: If the tokens are expected to align with texts (e.g. for semantic tokens and semantic-distilled acoustic tokens), metrics like phone discrminability [176], phone purity, phone-normalized mutual information [32] can be computed.
  • Invariance and robustness: If the tokens are expected to be invariant to perturbations, unit edit distance [38] can be considered.
There are several existing benchmarks on discrete speech tokens. Codec-SUPERB [^7] https://codecsuperb.com/ [186] evaluates both signal-level reconstruction metrics and downstream application performances of acoustic tokens. ESPnet-Codec [187] integrates multiple codecs into a unified training and evaluation framework (named VERSA [^8] https://github.com/shinjiwlab/versa) and extends evaluation to some generative tasks such as TTS and SVS. DASB [^9] https://github.com/speechbrain/benchmarks/tree/main/benchmarks/DASB [153] performs more downstream probing tasks, and includes generative tasks as well as semantic tokens. STAB [188] takes a different perspective that measures the invariance, robustness, compressibility, and vocabulary utilization of speech tokens. This emphasizes the application of discrete speech tokens in spoken language models instead of reconstruction. Expresso [189] introduces an expressive speech dataset and evaluates the resynthesis quality of HuBERT and EnCodec tokens, finding that HuBERT struggles to preserve source speech expressivity while EnCodec performs better.
@@ V-B Existing Analyses
There are theoretical or experimental analyses of the advantages of discrete speech tokens. Nguyen et al. [190] demonstrate by an encoder-only language model that discrete semantic SSL tokens are favorable for spoken language modeling, due to their removal of linguistically irrelevant information. Sicherman et al. [123] support this claim by showing a strong correlation of semantic units to phonemes but weaker to speakers. Abdullah et al. [191] refine the correlation between semantic SSL tokens and linguistic content to the “sub-phonemic” level instead of high-level phonemes due to contextual variability. Chang et al. [135] explores the use of WavLM tokens for end-to-end ASR together with deduplication and BPE. Although these tokens underperform continuous SSL features, they still show competitive performance. Similar findings are reported on contextual ASR [192], multilingual ASR [193], end-to-end speech translation, understanding [165], and more LLM-based semantic related tasks with discrete units as inputs [194].
The downside of speech tokens is also researched. Yeh et al. [124] investigates the information completeness (how much information is present) and accessibility (how prominent the information is) of HuBERT tokens. It suggests that VQ on HuBERT representations cannot achieve perfect disentanglement of speaker and phonetic content. EMO-Codec [195] compares the emotion recognition performance of the reconstructed speech from various acoustic tokens, showing that reconstruction of these tokens still sometimes degrades the emotion information. O’Reilly et al. [196] shows that neural audio codecs often lack stability after multiple rounds of encoding and decoding, i.e. not idempotent.
@@ V-C Reconstruction Analysis
To enable a fair comparison between acoustic and semantic tokens from a reconstruction perspective, we train a CTX-vec2wavα vocoder [28] for different semantic tokens on LibriTTS [197]. As semantic tokens usually lose much speaker timbre information, this vocoder supplements this information using continuous WavLM features extracted from the reference prompts. This approach enables semantic tokens to perform voice conversion (VC) by switching reference prompts conveniently. The training details follow [119]. We compute several metrics for reconstruction ability:
@@
We use LibriTTS testset-B proposed in [119] as the test set for evaluations. It contains 500 utterances from unseen speakers that sum up to about 1 hour. We use the original utterance to provide timbre information when necessary, i.e. TiCodec, FACodec, LSCodec and all semantic tokens [^12] Note that there is potential of information leakage in this situation. An alternative strategy is to use a reference prompt from the same speaker to provide timbre information for every utterance, then compute similarity between the generated utterance and the reference. However, this is rather “same-speaker conversion” instead of reconstruction, and the SECS values are not comparable to ordinary acoustic tokens that don’t require additional speaker input.. All evaluation metrics are computed on 16kHz waveform for a fair comparison, and reconstructed waveforms with higher sampling rates are downsampled before evaluation.
We take representative works in each token category. For conciseness, when there are multiple feasible configurations for an acoustic token, we choose one typical configuration that balances bitrate and performance. Note that different variants (especially on frame rate and number of quantizers) within the same model can lead to significant differences in reported metrics. For SSL models like wav2vec 2.0, HuBERT and WavLM, we take the official “Large” model variant. For wav2vec 2.0, we experiment with both its inner quantizer before the Transformer blocks and k-means clustering results on a specific Transformer layer.
The results are shown in Table III. It is evident that acoustic tokens designed only for reconstruction can achieve decent speech quality, but still far from perfect compared to traditional spectrogram-based vocoders because of higher compression rates. Retaining good speech intelligibility (i.e. low WER) becomes particularly challenging when the frame rate is low. Acoustic tokens with semantic distillation can also achieve strong reconstruction quality. Explicitly disentangled acoustic tokens may sacrifice some reconstruction performance metrics when the bitrate is extremely low (e.g. LSCodec). Semantic tokens generally struggle to achieve the same level of acoustic reconstruction as acoustic tokens, as evidenced by lower GPE, PESQ, and STOI scores. Notably, most semantic tokens included exhibit significant information loss in prosody as reflected by their GPE scores. However, their WER scores remain comparable to acoustic tokens, despite having much lower bitrates. This highlights the property that semantic tokens primarily retain content-related information rather than acoustic details.
@@ VI-A1 Motivation
Spoken language understanding (SLU) tasks, including automatic speech recognition (ASR), speech translation, slot filling, intent classification (IC), emotion recognition (ER) and others, aim to extract meaningful domain-specific information from speech. Most of these tasks follow a speech-in text-out framework, except S2ST which also involves speech generation. The adoption of discrete tokens in SLU offers some benefits. Discrete tokens may naturally exhibit a certain degree of robustness against noise, which can enhance the reliability of the system. Also, since semantic tokens often remove much speaker information, they enable downstream models to focus more effectively on content-related information, particularly on speaker-irrelevant tasks. On a broader scale, discrete tokens provide a promising approach to jointly learning speech understanding and generation in spoken language models.
As an alternative input to an SLU model instead of continuous features, discrete speech tokens are typically deduplicated or BPE-encoded, before feeding to the subsequent modules. Semantic tokens have been better explored than acoustic ones in this context.
@@ VI-A2 Speech Translation
Among the various SLU tasks, discrete speech tokens are mostly adopted in speech translation. Speech translation involves two primary tasks: speech-to-text translation (S2TT) and speech-to-speech translation (S2ST). Since semantic tokens are shown to correlate well with phonetics, they can serve as a universal pseudo-label for untranscribed languages. As a result, in low-resource settings, leveraging discrete tokens has demonstrated significant potential for improving S2TT performance [198].
Direct S2ST using discrete tokens has garnered more attention on the generation side, which will also be explained in Section VI-B. Early approaches primarily rely on extracting discrete tokens using VQ-VAE models, particularly for speech translation involving unwritten languages [199, 200]. With the growing adoption of SSL models, research in this area has expanded to include semantic tokens [201, 202]. Subsequent studies have focused on enhancing translation quality and efficiency through various methods, such as introducing two-pass architectures [203, 204], employing acoustic tokens [205, 206, 207], and developing non-autoregressive (NAR) frameworks [208]. These efforts collectively contribute to advancing the performance and applicability of discrete-token-based speech translation systems.
@@ VI-A3 Unified Speech Understanding
Since the SpeechPrompt series of work [209, 210, 211] first explored the SLU capabilities of the spoken language modeling framework through prompt tuning, discrete tokens have been explored for constructing spoken language models with SLU functionality. These models leverage discrete tokens to improve efficiency and adaptability across various tasks. AudioPaLM [212] extends the PaLM architecture [213] by incorporating audio embeddings and jointly modeling text and audio tokens, demonstrating abilities in ASR and S2TT by leveraging a shared vocabulary for both modalities. VioLA [7] adapts the VALL-E [6] framework by interleaving speech and text tokens and employing autoregressive modeling across both modalities. This design enables VioLA to handle various SLU tasks, including ASR, S2TT, and S2ST LauraGPT [214] also shows notable capabilities in generative tasks besides understanding, such as TTS and speech enhancement. These efforts highlight the potential of discrete tokens in enhancing the performance and versatility of universal SLU models.
@@ VI-A4 Limitations
Despite the advantages of reduced transmission cost and growing popularity in S2ST tasks, discrete tokens still underperform in many SLU tasks. Lots of studies on discrete token-based SLU models [215, 165, 216, 192, 193] only verify that discrete tokens can surpass continuous mel-spectrogram features in certain tasks such as ASR. Other continuous features, particularly SSL features, continue to provide superior performance especially when integrated with language models [194]. The majority of current LLM-based SLU models rely predominantly on continuous inputs, such as Whisper features [217, 218, 219, 220, 221, 222]. Moreover, the performance of discrete tokens in speaker-related tasks is generally inferior to that of continuous features by a large margin [215, 153]. Compared to continuous features, a significant limitation of discrete tokens for SLU is the inevitable information loss during the quantization process. Although acoustic tokens with sufficient decompression ability will mitigate such loss, they may also hinder the accessibility of semantic information which is crucial for SLU. Therefore, the full potential of leveraging discrete tokens for SLU remains largely untapped and warrants further exploration.
@@ VI-B1 Motivation
Discrete tokens have catalyzed a paradigm shift in speech generation, with TTS being the most representative application. In TTS systems, discrete tokens are usually used as intermediate features that bridge the acoustic model (text-to-token) and the vocoder or codec decoder (token-to-wav). The utilization of discrete speech tokens has driven the evolution of TTS models towards high-fidelity, zero-shot and controllable synthesis. There are two major advantages of applying discrete tokens in TTS:
  • Easier training objectives. Discrete tokens replace the original spectrogram-based regression task with a classification task [136]. As the discrete tokens take values from a finite set, predicting them is easier for acoustic models than continuous features which can be infinitely precise. In traditional TTS paradigms [223, 224], the vocoder takes a relatively easy task to convert frequency-domain features into waveforms, while the acoustic model transforms text into a largely different domain. Discrete tokens help to better balance the modeling burden between the two processes, thus makes the prediction task of the acoustic model more manageable.
  • Better use of decoder-only language models. Decoder-only language models have shown remarkable success in natural language generation. After discretization, speech can also be autoregressively generated under the same paradigm. This offers huge potential in leveraging the in-context learning capabilities of language models to achieve zero-shot controllable TTS [6]. Moreover, such TTS models can achieve considerable performance improvements after scaling the data and model size, leading to high-fidelity synthesis [112, 185].
@@ VI-B2 Autoregressive TTS
Autoregressively predicting the next VQ index of discrete speech tokens is first proposed in VQTTS [136], which is modeled by an LSTM conditioned on duration-upsampled phone representations and previous token outputs. A discrete token vocoder trained on a single-speaker corpus converts the tokens to waveforms with the assistance of handcrafted prosody features. This system uses vq-wav2vec tokens. VQTTS achieves state-of-the-art TTS quality at that time, and shows promising performance in speaker-adaptive TTS [180, 181, 182], expressive TTS [228] and TTS with low-bitrate tokens [229].
Subsequently, decoder-only TTS models using neural audio codecs have made tremendous success in zero-shot TTS starting from VALL-E [6]. VALL-E contains an autoregressive (AR) model and non-autoregressive (NAR) model. The AR model performs next-token prediction on the sequence of text and the first RVQ layer from EnCodec. The NAR model receives the text, all EnCodec codes from the speaker reference and the previous $k$ EnCodec RVQ levels to predict the codes at $k+1$-th level together. VALL-E operates under a concise design where the text and speaker reference are both considered “prompts” of a language model. It achieves remarkable zero-shot TTS performance when trained on 60k hours of speech. Later, methods have been proposed to improve generation robustness [230, 231, 232, 233, 234, 235], efficiency [236], style control [237, 238, 239], and to incorporate LLMs [240, 241].
Besides using an NAR model to predict the rest RVQ layers, alternate modeling strategies have been proposed. UniAudio [242] uses hierarachical Transformers to autoregressively predict fine-grained RVQ codes in each frame. Token interleaving patterns [243, 244] can further simplify the model architecture to only one AR Transformer with multiple prediction heads. Semantic tokens are also introduced to cooperate with acoustic codecs [14, 137, 241], which might decrease the modeling difficulty since they bridge the gap between texts and acoustics and usually require only a single token stream. Such simplification can provide better streaming ability by speech-text interleaving [245]. Numerous industry-level large-scale TTS systems have been produced in this autoregressive TTS paradigm, such as XTTS [246], BASE-TTS [112], Seed-TTS [185], CosyVoice [40, 162], Fish-Speech [247], etc.

### v2

@@ IV-F Speech Token Vocoders
Polyak et al. [143] first explores speech resynthesis from discrete speech units by a HifiGAN [170] augmented with discretized pitch units and speaker embedding inputs. The vec2wav vocoder in VQTTS [129] improves this vocoder by a Conformer [93] frontend module before HifiGAN generator. Later, CTX-vec2wav [111] proposes a position-agnostic cross-attention mechanism that effectively integrates timbre information from surrounding acoustic contexts without the need for pretrained speaker embeddings. This makes it more timbre-controllable and suitable for zero-shot TTS and VC [112]. Upon it, vec2wav 2.0 [171] further advances the timbre controllability by SSL timbre features and adaptive activations, demonstrating a strong VC performance.
It is also feasible to apply diffusion or flow matching algorithms in token vocoders [172, 173, 41]. There, the discrete tokens are treated as a condition for diffusion or flow matching to generate mel-spectrograms, and further converted to waveform by a pretrained mel vocoder. Compared to training a token-to-wav vocoder in an end-to-end fashion, training a token-to-mel model is more convenient and does not need adversarial training. To better control timbre, a mask strategy is introduced into the training process where the model only computes loss on the un-masked part of spectrograms [41]. During inference, spectrogram from speaker prompt conditions the generative process, which can be regarded as a form of “in-context learning”. However, this requires tokens to be extracted from reference prompts before synthesis. Also, inference efficiency may be compromised for better generation quality with multiple inference steps, and this method is only validated on massive amount of data currently.
@@ V-A Metrics and Existing Benchmarks
  • Signal-level reconstruction metrics: For reconstruction evaluations, signal-level metrics like PESQ, STOI, mel distance, GPE, etc. are often used.
  • Perceptual reconstruction metrics: Apart from signal-level metrics, there can also be perceptual evaluations of reconstruction performance. This includes intelligibility (often measured by WER, CER, or phone error rates), speaker similarity, subjective or proxy MOS tests, etc.
  • Performance on downstream tasks: Probing tasks can be used to measure the preservation or prominence of certain information in tokens, like ASR, ASV, emotion recognition, and spoken language modeling [167]. Note that this is different from perceptual reconstruction metrics since it operates directly on tokens. Performance in generative tasks like TTS and VC can also be evaluated.
  • Semantic/phonetic relevance: If the tokens are expected to align with texts (e.g. for semantic tokens and semantic-distilled acoustic tokens), metrics like phone discrminability [167], phone purity, and phone-normalized mutual information [33] can be computed.
  • Invariance and robustness: If the tokens are expected to be invariant to perturbations, unit edit distance [39] can be considered as a measurement.
There are several existing benchmarks on discrete speech tokens. Codec-SUPERB [^5] https://codecsuperb.com/ [174] evaluates both signal-level reconstruction metrics and downstream performances of acoustic tokens. ESPnet-Codec [175] integrates multiple codecs into a unified training and evaluation framework VERSA [176] and extends evaluation to some generative tasks such as TTS and SVS. DASB [^6] https://github.com/speechbrain/benchmarks/tree/main/benchmarks/DASB [144] performs more downstream probing tasks, and includes generative tasks as well as semantic tokens. STAB [177] takes a different perspective that measures the invariance, robustness, compressibility, and vocabulary utilization of speech tokens. This emphasizes the application in spoken language models instead of reconstruction.
@@ V-B Existing Analyses
There are several theoretical or experimental analyses of the advantages of discrete speech tokens. Nguyen et al. [178] demonstrates by an encoder-only language model that semantic SSL tokens are favorable for spoken language modeling, due to their removal of linguistically irrelevant information. Sicherman et al. [115] supports this claim by showing that semantic units have a strong correlation with phonemes, but a weaker correlation with speakers. Abdullah et al. [179] refines the correlation between semantic SSL tokens and linguistic content to the “sub-phonemic” level instead of high-level phonemes due to contextual variability. Chang et al. [128] explores the use of WavLM tokens for end-to-end ASR together with deduplication and BPE. Although these tokens underperform continuous SSL features, they still show competitive performance. Similar findings are reported on contextual ASR [180], multilingual ASR [181], end-to-end speech translation, understanding [156], and more LLM-based semantic-related tasks with discrete units as inputs [182]. Expresso [183] evaluates the resynthesis quality of HuBERT and EnCodec tokens on an expressive dataset, finding that HuBERT struggles to preserve source speech expressivity while EnCodec performs better.
The downside of speech tokens is also researched. Yeh et al. [116] suggests that VQ on HuBERT embeddings cannot achieve perfect disentanglement of speaker and phonetic content. EMO-Codec [184] shows that codec reconstruction still sometimes degrades the emotion information. O’Reilly et al. [185] shows that neural audio codecs often lack stability after repeated encoding and decoding, i.e. not idempotent.
@@ V-C Reconstruction Analysis
To enable a fair comparison between acoustic and semantic tokens from a reconstruction perspective, we train a CTX-vec2wavα vocoder [29] for different semantic tokens on LibriTTS [186]. This vocoder supplements the insufficient speaker timbre information in semantic tokens using continuous WavLM features extracted from the reference prompts. This approach enables semantic tokens to perform voice conversion (VC) by switching reference prompts conveniently. The training details follow [111]. We compute several metrics for reconstruction ability:
@@
We use LibriTTS testset-B  [111] as the test set for evaluations. It contains 500 utterances from unseen speakers that sum up to about 1 hour. We use the original utterance to provide timbre information when necessary, i.e. for TiCodec, FACodec, LSCodec and all semantic tokens [^9] Note that there is potential of information leakage in this situation. An alternative strategy is to use a reference prompt from the same speaker to provide timbre information, then compute similarity between the generated utterance and the reference. However, this is rather “same-speaker conversion” instead of reconstruction, and the SECS values are not comparable to ordinary acoustic tokens that don’t require additional speaker inputs.. All evaluation metrics are computed on 16kHz waveform for a fair comparison, and reconstructed waveforms with higher sampling rates are downsampled before evaluation.
We take representative works in each token category. When there are multiple feasible configurations for a model, we choose one typical configuration that balances bitrate and performance. Note that different variants (especially on frame rate and number of quantizers) within the same model can lead to significant differences in reported metrics. For SSL models like wav2vec 2.0, HuBERT and WavLM, we take the official “Large” model variant. For wav2vec 2.0, we experiment with both its inner quantizer before the Transformer blocks and k-means clustering results on a specific Transformer layer.
The results are shown in Table III. It is evident that acoustic tokens designed only for reconstruction can achieve decent speech quality, but still far from the state-of-the-art spectrogram-based vocoders because of higher compression rates. Retaining good speech intelligibility (i.e. low WER) becomes particularly challenging when the frame rate is low. Acoustic tokens with semantic distillation can also achieve strong reconstruction quality. Explicitly disentangled acoustic tokens may sacrifice some reconstruction performance metrics when the bitrate is extremely low. Semantic tokens generally struggle to achieve the same level of acoustic reconstruction as acoustic tokens, as evidenced by lower GPE, PESQ, and STOI scores. Notably, most semantic tokens included exhibit significant information loss in prosody as reflected by their GPE scores. However, their WER scores remain comparable to acoustic tokens, despite having much lower bitrates. This highlights the property that semantic tokens primarily retain content-related information rather than acoustic details.
@@ VI-A1 Motivation
Spoken language understanding (SLU) tasks, including automatic speech recognition (ASR), speech translation, intent classification and others, aim to extract meaningful domain-specific information from speech. Most SLU tasks follow a speech-in text-out pipeline, except S2ST which also involves speech generation. The adoption of discrete tokens in SLU offers some benefits. Discrete tokens may naturally exhibit some invariance against noise and speaker information, particularly semantic tokens, which can make downstream models to focus more effectively on content-related information in some tasks. On a broader scale, discrete tokens provide a promising approach to unifying speech understanding and generation in spoken language models.
As an alternative input to an SLU model instead of continuous features, discrete speech tokens are typically deduplicated or BPE-encoded before subsequent modules. Semantic tokens have been better explored than acoustic ones in this context.
@@ VI-A2 Speech Translation
Among the various SLU tasks, discrete speech tokens are mostly adopted in speech translation, including speech-to-text translation (S2TT) and speech-to-speech translation (S2ST). Since semantic tokens correlate well with phonetics, they can serve as universal pseudo-labels for untranscribed languages, useful for S2TT in low-resource settings [187]. Direct S2ST using discrete tokens has garnered more attention on the generation side (Section VI-B). Early approaches primarily rely on extracting discrete tokens using VQ-VAEs, particularly for unwritten languages [188, 189]. Recent researches in this area include employing semantic tokens [190, 191, 192], acoustic tokens [193, 194, 195], two-pass architectures [196, 197], and non-autoregressive frameworks [198]. These efforts collectively contribute to advancing the performance and applicability of discrete token-based speech translation systems.
@@ VI-A3 Unified Speech Understanding
Discrete tokens provide opportunity to construct unified and adaptable spoken language models with various SLU functionalities. Efforts include task identifies [7], prompt tuning [199, 200, 201, 192], shared audio and text vocabulary [202], and combining continuous and discrete representations [203]. These efforts highlight the potential of discrete tokens in enhancing the performance and versatility of universal SLU models.
@@ VI-A4 Limitations
Despite the advantages and growing popularity in S2ST tasks, discrete tokens still underperform in many SLU tasks. Lots of SLU studies [204, 156, 205, 180, 181] only verify that discrete tokens can surpass traditional frequency-domain features in certain tasks such as ASR. Continuous SSL features continue to have superior performance [182]. The majority of current LLM-based SLU models rely predominantly on continuous inputs, such as Whisper features [206, 207, 208, 209, 210, 211]. Moreover, the performance of discrete tokens in speaker-related tasks is generally much inferior to that of continuous features [204, 144]. A significant limitation of discrete tokens for SLU is the inevitable information loss during the quantization process. Mitigating such loss with more VQ codebooks may hinder the accessibility of semantic information crucial for SLU as well. Therefore, the full potential of leveraging discrete tokens for SLU remains largely untapped and warrants further exploration.
@@ VI-B1 Motivation
Discrete tokens have catalyzed a paradigm shift in speech generation, with TTS being the most representative application. In TTS systems, discrete tokens are usually used as intermediate features that bridge the acoustic model (text-to-token) and the vocoder or codec decoder (token-to-wav). There are two major advantages of applying discrete tokens in TTS:
  • Easier training objectives. Discrete tokens replace the original spectrogram-based regression task with a classification task [129], which can be much easier. This also offers a better balance between acoustic models and vocoders, since texts are closer to discrete speech tokens than frequency-domain features.
  • Better use of decoder-only language models. Decoder-only language models have shown remarkable success in natural language generation. After discretization, speech can also be autoregressively generated under the same paradigm. This offers huge potential in leveraging the in-context learning and scaling capabilities of language models to achieve zero-shot high-fidelity TTS [6].
@@ VI-B2 Autoregressive TTS
Autoregressively predicting the next VQ index of discrete speech tokens is first proposed in VQTTS [129], which uses an LSTM conditioned on Transformer representations to generate vq-wav2vec [31] semantic tokens. A discrete token vocoder converts the tokens to waveforms with the assistance of handcrafted prosody features. VQTTS achieves state-of-the-art TTS quality at that time, and shows promising performance in speaker-adaptive TTS [215, 216, 217] and expressive TTS [218].
Subsequently, decoder-only TTS models using neural audio codecs have made tremendous success in zero-shot TTS starting from VALL-E [6]. VALL-E contains an autoregressive (AR) model and non-autoregressive (NAR) model, both of which generate EnCodec [63] RVQ tokens. The AR model performs next-token prediction on the first RVQ layer conditioned on text. The NAR model predicts the $n+1$-th RVQ tokens given the text, all EnCodec tokens from the speaker reference, and the previous $n$ RVQ layers. VALL-E employs a concise design in which text and speaker references serve as “prompts” for a language model. It achieves remarkable zero-shot TTS performance when trained on 60k hours of speech. Later, methods have been proposed to improve generation robustness [219, 220, 221, 222, 223, 224], efficiency [225], style control [226, 227, 228], and to incorporate LLMs [229, 230].
Besides using an NAR model to predict the rest RVQ layers, alternate modeling strategies have been proposed, such as hierarchical modeling [231] and token interleaving patterns [232, 233]. Semantic tokens are also introduced to cooperate with acoustic codecs [15, 130, 230, 234], which might decrease the modeling difficulty since they bridge the gap between texts and acoustics and usually require only a single token stream. Numerous industry-level large-scale TTS systems have been produced in this autoregressive TTS paradigm, such as XTTS [235], BASE-TTS [236], Seed-TTS [173], CosyVoice [41, 153], Fish-Speech [237], etc.